# High-Quality Axions from Higher-Form Symmetries in Extra Dimensions

### v1

@@ V Gravitational Considerations
The axion quality problem is often framed in terms of global symmetry-breaking effects from quantum gravity. Although the effects we have considered in our 5D toy model are all field-theoretic in nature, they may be used to infer the size of expected corrections to the extra-dimensional axion potential from quantum gravity via various forms of the Weak Gravity Conjecture (WGC) [67, 68, 26, 28]. While many of the WGC-induced connections between the extra-dimensional axion quality problem and quantum gravity expectations have been drawn elsewhere, they dovetail nicely with the present classification.
@@ VI Conclusion and Outlook
Additionally, we have discussed how the generalized Weak Gravity Conjectures constrain the parameters involved in electric one-form symmetry breaking in the 5D actions considered. It would be interesting to study more detailed quantum gravity inspired models with a more complex topology from the higher-form symmetry breaking point of view. Moreover, the one-form electric symmetry breaking discussed in this work is interpretable in a dual frame where it is the magnetic higher-form symmetry of the dual axion Kalb-Ramond field that is broken. We leave the study of these more complicated models and the magnetic dual for future endeavours.

### v2

@@ V Gravitational Considerations
The axion quality problem is often framed in terms of global symmetry-breaking effects from quantum gravity. Although the effects we have considered in our 5D toy model are all field-theoretic in nature, they may be used to infer the size of expected corrections to the extra-dimensional axion potential from quantum gravity via various forms of the Weak Gravity Conjecture (WGC) [68, 69, 26, 28]. While many of the WGC-induced connections between the extra-dimensional axion quality problem and quantum gravity expectations have been drawn elsewhere, they dovetail nicely with the present classification. Similarly, the Completeness Hypothesis imposes requirements on the spectrum of states in our theory, and these requirements in turn determine how the electric and magnetic higher-form symmetries associated with the axion gauge field are broken.
@@
V.4 Completeness Hypothesis
In its most basic form, the Completeness Hypothesis postulates that all charges consistent with Dirac quantization must be realized by physical states in a gauge theory that is consistently coupled to gravity [88, 89]. For the extra-dimensional axion of the form discussed in this work, the Completeness Hypothesis demands the existence of unit electric and unit magnetic charge states (but unlike the weak gravity conjecture, no restriction is placed on the masses of these states). Thus, both the $U(1)_{e}^{(1)}$ and $U(1)_{m}^{(2)}$ symmetries must be broken completely, without the possibility of a surviving remnant discrete symmetry. This generally leads to $2\pi$ periodic potentials of the form discussed in Sec. III.1 and IV  [^8] We note that the Completeness Hypothesis has been studied and generalized in the context of higher-form gauging and non-invertible symmetries, where the notion of completeness of the spectrum can be shown to be equivalent to the endability of extended operators [96]..
@@ VI Conclusion and Outlook
Additionally, we have discussed how generalized weak gravity conjectures and the Completeness Hypothesis constrain the parameters involved in electric one-form symmetry breaking in the 5D actions considered. It would be interesting to study more detailed quantum gravity inspired models with a more complex topology from the higher-form symmetry breaking point of view. Moreover, the one-form electric symmetry breaking discussed in this work is interpretable in a dual frame where it is the magnetic higher-form symmetry of the dual axion Kalb-Ramond field that is broken. We leave the study of these more complicated models and the magnetic dual for future endeavors.